1. a data set includes 83 observations. How many classes would you recommend to develop a frequency distribution?

2. a data set consists of 145 observations ranging from 56 to 490. What size class range would you recommend?
Answer #1

Solution:

1. The number of classrs will be 7

Explanation:

in this case, if k = 6 then 2^k = 2^6 = 64 which is the less than n = 83. so,K = 6 will not work.if k = 7, then = 2^7 = 128 which is greater than n = 83.so number of classes will be 7.

2.
Class interval = 50

in the case, n = 145. if K = 5. then 2^k = 2^7 = 128 which is less than n = 125.So k = 7 will not work.if k = 8,then 2^k = 2^8 = 256 which is greater than n = 145. so number of classes will be 8.


=> i > = (490-56)/8 = 54.25